OpenStudy (anonymous):

Find the length and width (with W < or equal to L ) of the rectangle with perimeter 72 that has maximum area, and then find the maximum area.

OpenStudy (anonymous):

2W+2L=72 W+L=36 W=36-L A=LW=L(36-L) you want to find the vertex of the parabola A=L(36-L) the x-coordinate is your length, y-coordinate is the maximum area

OpenStudy (campbell_st):

the easy way is to use a spreadsheet the sum of length and width is 36 so start with a spreadsheet with columns Width Length Area 1 35 35 2 34 78 3 33 99 repeat the process graph the points and you'll see a concave down parabola. the max area occurs when l = w = 18 or you can use differential calculus

OpenStudy (anonymous):

Thank you! :)